OpenStudy (anonymous):

rewrite sqrt(4x^2-9) in the form k*sqrt(x^2-a^2)

OpenStudy (anonymous):

do u know that \[(a -b)*(a+b) = a^2 - b^2 \] here can you see any relation?

OpenStudy (anonymous):

surjithayer is typing a reply… :P LOL just kidding :P :D

OpenStudy (anonymous):

\[\sqrt{4x ^{2}-9}=2\sqrt{x ^{2}-\frac{ 9 }{4 }}=2\sqrt{x ^{2}-\left( \frac{ 3 }{ 2 } \right)^{2}}\]